OpenStudy (anonymous):

Simplify 5 log2 k – 8 log2 m + 10 log2 n

OpenStudy (anonymous):

what have you tried ?

jhonyy9 (jhonyy9):

so i think this will be right

jhonyy9 (jhonyy9):

\[\log2(k ^{5}/(m ^{8}*n ^{10}))\]

OpenStudy (anonymous):

5.log2 k -8.log2 m +10.log2 n= log2 (K)^5 -log2 (m)^8 + log2 (n)^10= log2{[(k)^5*(n)^10]/[(m)^8]}